According to the Bohr model of the atom, an electron in the ground state A. B. D. 41. radiates electromagnetic energy continuously emits only spectral lines can jump to another orbit if given enough energy C. remains there forever A hydrogen atom is said to be in its ground state when its electron A. isat rest C. is in the lowest energy level D has escaped from the atom An atom emits a photon when one of its orbital electrons A. jumps from a higher to a lower energy level B. jumps from a lower to a higher energy level C is removed by the photoelectric effect 42. B. is inside the nucleus 43. D. is struck by an X-ray 44. Light of wavelength 5 x 107m consists of photons whose energy is 45. The frequency of an X-ray photon which has an energy of 6.6 x 10-15 J is 46. The de Broglie wavelength of an electron whose speed is 10* m/s is 47. The speed of an electron whose de Broglie wavelength is 1010 m is A. 1.1 x 10 B 1.3 x 10 C 4x 109J D. 1.7x 101 J A. 4.4 x 10Hz B. 10Hz C.10 Hz A 5.9 x 106 m D. 10Hz B. 1.5x 10-19m C 7.3 x 10 12m D. 1.4x 10 m A 6.6x 102 m/s B. 3.8 x 10m/s. C. 7.3 x 100m/s D. 101 m/s The fluid at the top of a container is A. under less pressure than the fluid at the bottom B. under the same pressure as the fluid at the bottom C. under more pressure than the fluid at the bottom D.any of these, depending upon the circumstances A cake of soap placed in a bathtub of water sinks. The buoyant force on the soap A. is 0 ?.is less than its weight A 200-g piston rests on a sample of a gas in a cylinder whose cross-sectional area is IS m2. The pressure the piston exerts on the gas is A. 0.13 Pa B. 1.3 Pa C. 13Pa 48. 49. B, is equal to its weight D.?s more than its weight s0. D, 7.5 Pa ,

media%2F3f6%2F3f625f9c-4077-4d87-a03d-bc
Answer #1

40) density = mass / Volume

volume = length * breadth * height

V = lbh = 10*6*3 = 180 m^3

Density = 400 / 180 = 2.22 kg / m^3

so option A) 2.22 kg /m^3 is the correct answer